### v1

@@ 3.3 Main Beam Fitting
We fit a 1d Gaussian function to the flight data in TelX to determine four parameters, as shown in Figure 5:
$P(\nu,\phi)=Ae^{\frac{(\phi-\phi_{0})^{2}}{2\sigma^{2}}}+B$ (1)
where $A$ is the amplitude, $B$ is the background level, $\phi_{0}$ is the centroid location in TelX, and $2.355*\sigma$ is the full-width-half-max (FWHM). Each combination of feed index, frequency, and polarization is analyzed separately for the drone and holography datasets for CHIME cylinder C.
Figure 5: CHIME data acquired during the drone flight. The power measured by feed 190 (corresponding to TelY coordinate -0.59 and declination 13.15) is shown at three frequencies in both polarizations (XX top, YY bottom). The power has been normalized by the amplitude of the best-fit 1d Gaussian from Equation 1 (dotted line). The beam narrows with increasing frequency, as expected, and frequency-dependent sidelobe structure beyond the main beam shape is also apparent in the data.
@@ 3.4 Altitude Estimation
While exploring the evolution of the FWHM parameter in frequency and declination space we were able to determine the drone flight altitude with only a mild prior on the drone altitude from the drone sensors. This best-fit altitude differed from the altitude reported by the RTK system by a constant offset of $\sim 5$ m.
@@
In Figure 6 we present the FWHM in each polarization as a function of frequency and TelY for all three data sources. The FWHM parameter for each data product has been normalized by the median value in each frequency bin. The solar data spans the shown coordinate space, and is present in the background of each plot in the figure. The holography data is overlaid with the same normalization at the associated source declinations, neatly fitting within the ripple pattern present in the solar FWHM results. The best-fit FWHM results for the drone data are overlaid at the associated TelY [.62,.49]absent.62.49$\in$[-.62,-.49]∈ [ - .62 , - .49 ], and fall within the median-valued contour for the solar data at the associated declination.
Figure 6: Best-fit FWHM as a function of frequency and altitude for the XX (top) and YY (bottom) polarizations. In each panel, the Solar data spans the coordinate space, the Drone data is overlaid for TelY within the range of (-0.62,-0.48), and the Holography measurements from Virgo A are plotted as a band of points at TelY0.6similar-toabsent0.6$\sim$-0.6∼ - 0.6. The colors are normalized to the median value within each frequency bin, and a contour at the median value has been traced over the solar data to enhance the appearance of the 30MHz ripple feature. The evolution of the FWHM parameter follows the same contour for all three data sets, and shows the characteristic 30MHz frequency ripple moving predictably as the altitude/declination increases.
@@
Using an altitude value of 307.6 m (311.6 m above the image plane) from the drone’s internal altimeter and RTK sensors, the drone FWHM results appeared offset by TelY $\approx-0.005$ relative to the solar FWHM pattern. We estimated that a $\sim 5$ m offset in flight altitude would account for this shift in TelY, suggesting the flight altitude was instead 312.6 m (316.6 m above the image plane). To further quantify this offset, we use a Pearson R (PR) correlation coefficient computation as follows. Our geometric model of the drone flight prescribes a TelY value to each telescope feed, defined by the position and altitude of the drone. We can independently determine a mapping between telescope feed and TelY by computing the correlation between the drone and solar FWHM curves. Specifically, for each telescope feed in the drone data we determine the TelY value in the solar data that maximizes the PR correlation coefficient. This associates each feed in the drone data with a best-fit TelY value. This is computed for each feed independently and we do not impose any requirements (e.g., that the FWHM pattern of all feeds taken together must match the solar data with a given geometric model).
@@
The PR coefficients are shown in Figure 7, associating each of the 256 feeds used in the drone measurement (x-axis) with a Solar TelY value (y-axis). The results are roughly consistent with a constant altitude offset, a flight height of 315.1 m above the ground (319.1 m above the image plane), indicated in the figure. Figure 7 also shows that the PR coefficients give a unique mapping between TelY and feed location for the YY polarization; however for the XX polarization, additional offsets also correlate strongly, although still has a clear maximum at the same location as the YY solution. As we’ll show below, we suspect this is because the FWHMs for the XX polarization have more RFI-flagged frequencies and additional substructure in the FWHM pattern beyond the 30 MHz ripple. Figure 7 also shows the maximum PR points and the varying altitude curves after subtracting the TelY per feed mapping prescribed by the geometric model. This analysis suggests the magnitude of the global altitude offset is +7.5 m relative to the altitude reported by the drone. After this correction, a clear linear trend remains between TelY and chime feed. This may indicate there are unaccounted-for offsets (e.g. in the latitude and longitude sensors for the drone), residual effects from the different processing between the data sets (drone data is per-feed, while solar data is formed from short baselines), or effects from the near-field evolution the measurement). In addition, the residual sawtooth pattern is reminscent of the centroid offsets per feed (see Section 3.3) which were not accounted for in this analysis.
@@
Figure 7: Pearson-R (PR) correlation coefficients calculated between drone data and solar data for each polarization (top XX, middle YY) of the 256 CHIME feeds mapped by the drone. The maximum value in each column (white dots) associates a CHIME feed with a TelY bin in the solar data, restricted to lie within bounds (red lines) to eliminate spurious outliers. Also shown are colored arcs indicating various drone flight altitudes above the ground level. The maximum PR coefficient between the drone and Virgo A holography data occurs at feed 210 (yellow star) and falls along the TelY and feed mapping obtained from correlations with solar data. The bottom panel shows the colored arcs and maximum PR correlation coefficients (white and purple dots) for both polarizations after subtracting the TelY and feed mapping prescribed by the geometric model.
@@ 4.1 Centroid and FWHM Measurements
From the best-fit 1d Gaussian, we obtain centroid and FWHM parameters as described in 3.3. Both polarizations are measured independently in subsequent transits at the same zenith angle. Figure 8 shows the best-fit centroid values from the drone and holography data and compares them to photographic measurements performed along the CHIME feedline to measure feed displacement along the focal line [41] Because the solar data does not include any feed-level data, it is not represented in Figure 8. Although the drone measurements are in the near-field, the centroid values and trend along the focal line from the drone agree with the two far-field measurements and follow the feed displacements. The centroid trend is known, and was previously published in [1].
@@
Figure 9: Measured FWHM from Drone, Holography (both for a single feed and an average of baselines $B\leq 10m$ to form a data set equivalent to the solar data), and Solar data sets as a function of TelX. These curves can be directly compared because they all originate from a TelY of -0.6, where all three data sets overlap (see Figure 4). The holography and drone measurements are most comparable, because they can be separated into individual feeds. Conversely, the solar measurements are comprised of an average of all intercylinder baselines under 10m. To examine the divergence of the YY Polarization drone FWHM curve from the solar and holography datasets, we investigate transit data in the frequency band (676.2-695.7 MHz) within the purple region in Figure 10.
The FWHM parameter is compared for all data sets at TelY $\approx-0.6$ as a function of frequency in Figure 9, and in particular show the solar data (formed from combining $B\leq 10m$), the drone and holography data from a single feed, and the holography data combined to form a data set equivalent to the solar data ($B\leq 10m$). The characteristic 30 MHz ripple feature is aligned for both polarizations at all frequencies, as one would expect from Figure 6. The amplitude of the ripple feature is is generally comparable between all data sets for the XX polarization, though the solar FWHM is wider than the holography and drone data sets throughout the frequency band. For the YY polarization, we find that the combined-feed holography data matches its solar counterpart very well, while the ripple amplitude is slightly higher for the single-feed holography data compared to the combined-feed holography data. This suggests that feed-to-feed variations may influence the amplitude of the ripple feature. We also find the peak FWHM amplitude found in the drone data is typically significantly higher than the other data sets for YY polarization, which prompted further investigation.
Figure 10 contains drone transit measurements for frequencies within 676.2-695.7 MHz. This frequency band follows the evolution of the 30 MHz ripple feature in the FWHM$(f)$ curve from a maximum to a minimum, indicated by purple region in Figure 9. In both polarizations, we expect to find a well defined main beam at TelX=0.0 that ends at a first null around ±$\pm 0.04\leq\textrm{TelX}\leq\pm 0.05$). Wherever the 30 MHz ripple feature reaches a local minimum, (e.g. $f=695$ MHz plotted in black in Figure 10,) the main beam and sidelobes are discrete and the best-fit Gaussian FWHM values agree with the solar and holography datasets as expected for both the XX (top panel) and YY (bottom panel) polarizations. Conversely, wherever the 30 MHz ripple reaches a local maximum, the primary beam and first sidelobe are not separated by a null in the drone transit data. Instead, the main beam and first sidelobe are combined and the Gaussian fit finds a FWHM that is wider than expected. Examples can be found in both the XX polarization (e.g. $f=685.0$ MHz plotted in pink in the top panel of Figure 10) and YY polarization (e.g. $f=677.5$ MHz plotted in yellow in the bottom panel of Figure 10).
To demonstrate this behavior, we compare both polarizations at $f=677.5$ MHz (plotted in yellow). The XX polarization data clearly recovers the main beam and two well-defined sidelobes. The YY polarization data only recovers the second sidelobe (centered at $\textrm{TelX}\approx\pm 0.1$) while the first sidelobe (centered at $\textrm{TelX}\approx\pm 0.05$ in XX) been been incorporated into the main beam. This behavior evolves smoothly in both polarizations as we progress along the gradient of plotted frequencies. The XX polarization recovers the expected first null on either side of the main beam at the highest and lowest plotted frequencies, but intermediate frequencies show the first sidelobe joining the main beam (particularly for TelX¡0). The evolution for the YY polarization is similar but not identical—as the plotted frequency decreases, the sharpness of the null diminishes, eroding the distinction between the main beam and the first sidelobe.
Figure 10: First eigenmode data from autocorrelation measurements of the drone transiting at TelY $\approx-0.6$ (feed 210) across CHIME cylinder C. The data shown are not normalized or corrected with the best-fit centroid. The frequency band shown here is reversed from the frequency axis of Figure 9. The colored curves thus evolve along the 30 MHz ripple feature in the FWHM$(f)$ curve starting from a minima at 695.7 MHz (black), where the FWHM$(f)$ values are in agreement with the solar and holography datasets, to a maxima at 676.2 MHz (yellow), where the divergence is highest. In the logarithmic scale shown in Figure 10 the amplitude appears compressed, but in linear units the amplitude does not appear compressed.
4.2 Trends with Frequency and TelY
The solar and drone data can be compared in the TelX vs. frequency and TelX vs. TelY subspaces to further investigate beam patterns. The drone data is taken from the first eigenmode (see Section 2.2.3) to increase the signal to noise. As previously described, the solar data is an average of all intercylinder baselines $B\leq 10m$. All data has been corrected for the frequency-dependent centroid deviation measured in the Gaussian fits.
First, we compare the drone and solar data as a function of TelX and frequency in Figure 11. For this comparison we select a single TelY index for the solar transit data and a single CHIME feed index for the drone data (both corresponding to $TelY=-0.600$, feed 210). The amplitude of the drone and solar data have been normalized in each frequency with the amplitude of the associated best-fit gaussian. These waterfall plots show similar frequency dependent variations in width in the main beam and sidelobe amplitudes and positions. Specifically, the width decreases with increasing frequency, and is modulated by a 30 MHz ripple aligned between the two data sets (as would be expected from Figure 9). Similarly, the XX polarization has a wider overall beam than the YY polarization, also as expected from the CHIME feed illumination pattern[42, 1]. The drone data has a more pronounced ripple in the XX polarization, whereas the ripple is attenuated in the solar data due to averaging over many baselines (supported by the lower ripple amplitude as described in Section 2.3.3).
For both polarizations, the main beam and sidelobes have similar TelY dependence in both the drone and solar datasets. A notable difference between these datasets is the high signal-to-noise and continuity present in the solar data from coadding in baseline and time. Conversely, each of the TelY values in the drone data comes from the same 3.4 min drone flight, and does not benefit from coadding or averaging, however signal-to-noise improvements are possible with modest increases in flight time.
Comparing the solar and drone data for the YY polarization, the following similarities are apparent: The main beam expands and shrinks at the same values of TelY (expands at TelY=-0.56, two clear ‘waists’ at TelY $\sim$-0.53 and -0.6). The sidelobe levels are higher where the beam widens (TelY=-0.56) and lower at TelY values where the main beam shrinks to a waist. The third order sidelobe is maximal at TelX=±$\pm 0.125$ for the plotted range of TelY, appearing as a continuous pink line.
@@ 4.2 Trends with Frequency and TelY
Turning to the XX polarization, we observe the same similarities across the drone and solar data. The main beam broadens at TelY=-0.55, and shrinks to a waist at TelY=-0.52 and TelY=-0.59 in both data sets. Unlike in polarization YY, the sidelobe levels for polarization XX evolve in TelX and TelY simultaneously—shown by symmetric sloped contours that extend away from waist in the main beam (e.g. starting at TelY=-0.59, and expanding in TelX as TelY decreases to -0.62.)
@@
Figure 11: Waterfall plots of telescope data obtained during Drone measurements (Right Column, single feed [210] in either polarization) and Solar measurements (Left Column, average of intercylinder baselines ($B\leq 10m$) in either polarization). Similar widths and features are present in both data sets, and the beamwidths and sidelobe features are comparable. Each frequency is normalized independently to a Gaussian fit peak value to remove the uncalibrated frequency dependence of the drone transmitter.
@@
Figure 12: Waterfall plots of telescope data obtained during Solar measurements (average of intercylinder baselines $B\leq 10$ for a single frequency bin (center 690MHz) in XX (Left Top) and YY (Left Bottom) polarizations) and Drone measurements (all feeds for a single frequency bin (center 690 MHz) in XX (Right Top) and YY (Right Bottom) polarizations).

### v2

@@ 3.3 Main Beam Fitting
We perform an unweighted least-squares fit of a 1d Gaussian function to the flight data in TelX to determine four parameters, as shown in Figure 5. For a single feed and frequency, the power $P$ as a function of TelX received as a source transits at a fixed TelY is:
$P(\mathrm{TelX})=Ae^{-\frac{(\mathrm{TelX}-\mathrm{TelX_{0}})^{2}}{2\sigma^{2}}}+B$ (1)
where $A$ is the amplitude, $B$ is the background level, $\mathrm{TelX_{0}}$ is the centroid location in TelX, and $2.355*\sigma$ is the FWHM. Each combination of feed index, frequency, and polarization is analyzed separately for the drone and holography datasets for CHIME cylinder C.
Figure 5: CHIME data (dotted lines) acquired during the drone flight. The power measured by feed 190 (corresponding to TelY coordinate -0.59 and declination 13.15) is shown at three frequencies in both polarizations ($X$ top, $Y$ bottom). The power has been normalized by the amplitude of the best-fit 1d Gaussian from Equation 1 (solid line). The beam narrows with increasing frequency, as expected, and frequency-dependent sidelobe structure beyond the main beam shape is also apparent in the data.
@@ 3.4 Altitude Estimation
In Figure 6 we present the FWHM in each polarization as a function of frequency and TelY for all three data sources. The FWHM parameter for solar and drone data at each frequency has been normalized by the median value across the TelY axis. The solar data spans the shown coordinate space, and is present in the background of each plot in the figure. The holography data (normalized with the median across the frequency axis instead) is overlaid at TelY=-0.6, neatly fitting within the ripple pattern present in the solar FWHM results. The FWHM results for the drone data, after accounting for the altitude estimation described in this section, are overlaid at the associated TelY $\in[-.62,-.49]$, and fall within the median-valued contour for the solar data at the associated declination.
Due to well-known multi-path effects within CHIME, the FWHM is a strong function of both TelY and frequency [31]. This can be seen in the background of Figure 6, which shows FWHM values derived from solar data as a function of frequency and TelY. The solar data clearly has curving stripes with a 30MHz frequency ripple (associated with the distance between the telescope vertex and focus) and the peaks of the ripples shift with location along the North-South axis (TelY). We can include the drone FWHM on these plots, and determine if our mapping between feed number and solar TelY values is correct. Because we are using the ripple pattern to assess the mapping between TelY and feed, we normalize by the median FWHM value in each frequency bin. Because the actual FWHM values do not always agree (see more detail in Section 4), this allows us to match using the contrast between the peaks and troughs, instead of the values themselves. While exploring the evolution of the FWHM parameter in frequency and TelY space we were able to determine the drone flight altitude with only a mild prior on the drone altitude from the drone sensors. This best-fit altitude differed from the altitude reported by the RTK system by a constant offset of $\sim 5$ m.
[…]
@@
Using an altitude value of 307.6 m above ground (311.6 m above the image plane) from the drone’s internal altimeter and RTK sensors, the drone FWHM results appeared offset by TelY $\approx-0.005$ relative to the solar FWHM pattern. We estimated that an $\sim 5$ m offset in flight altitude would account for this shift in TelY, suggesting the flight altitude was instead 312.6 m (316.6 m above the image plane). To find a more precise value of this offset, we use a Pearson R (PR) [^5] Pearson R correlation coefficient computation as follows. For a given feed, the drone position and altitude define an angle which we convert to TelY, as described in Section 3.1. Plotting data from all feeds produces a swath of FWHM values as a function of TelY and frequency, as shown in Figure 6. We can vary the altitude of the drone, which shifts the swath up or down in TelY, and compare the resulting FWHM values to the solar data in the same swath. If the swath in TelY for a given altitude is incorrect, the peaks and troughs in the drone data won’t align with the peaks and troughs in the solar data. We vary the drone altitude and compute the PR correlation coefficient between the drone and solar FWHM values. The altitude value which maximizes the PR coefficient provides a more precise estimate of the flight altitude than is given by the drone sensors. This is computed for each feed independently and we do not impose any requirements (e.g., that the FWHM pattern of all feeds taken together must match the solar data with a given geometric model).
@@
The PR coefficients are shown in the upper two panels of Figure 7, associating each of the 256 feeds used in the drone measurement (x-axis) with a Solar TelY value (y-axis). The results are roughly consistent with a constant altitude offset, a flight height of 315.1 m above the ground (319.1 m above the image plane), indicated by the black dashed lines in Figure 7. The two upper panels in Figure 7 also show that the PR coefficients give a unique mapping between TelY and feed location for the YY polarization; however for the XX polarization, additional offsets also correlate strongly, although still has a clear maximum at the same location as the YY solution. As we show below, we suspect this is because the FWHMs for the XX polarization have more RFI-flagged frequencies and additional substructure in the FWHM pattern beyond the 30 MHz ripple. This analysis suggests the magnitude of the global altitude offset is +7.5 m relative to the altitude reported by the drone.
Similarly, we compute the PR correlation between the drone and holography FWHM values across all frequencies for each feed. The maximum PR value thus represents which feed best matches the holography data from VirgoA (at TelY=-0.6). We find that the maximum value occurs at feed 210, which is indicated with a star in Figure 7. As expected, this independent determination of the feed to TelY relationship aligns with the mapping obtained from the drone and solar PR comparisons.
The lower panel of Figure 7 shows the maximum PR points after subtracting the TelY per feed mapping prescribed by the geometric model. After this subtraction, a clear linear trend remains between TelY and chime feed, reminiscent of the centroid offsets per feed (see Section 3.3) which were not accounted for in this analysis. In addition, a residual sawtooth pattern is present spaced roughly every 64 feeds. Because solar data combines measurements from all cylinders while the drone data is generated from a single cylinder we would expect differences between the two data sets due to cylinder-dependent effects.
@@
Figure 7: Pearson-R (PR) correlation coefficients calculated between drone data and solar data for each polarization (top $X$, middle $Y$) of the 256 CHIME feeds mapped by the drone. The maximum value in each column (color dots) associates a CHIME feed with a TelY bin in the solar data. Lines indicate the best-fit altitude (319.1 m). The maximum PR coefficient between the drone and Virgo A holography data occurs at feed 210 (yellow star) and falls along the TelY and feed mapping obtained from correlations with solar data. The bottom panel shows the and maximum PR correlation coefficients (blue and purple dots) for both polarizations after subtracting the TelY and feed mapping prescribed by the geometric model. An additional line is shown for the altitude given by the drone sensors (311.6 m) in the bottom panel.
@@ 4.1 Centroid and FWHM Measurements
From the best-fit 1d Gaussian (Equation 1), we obtain centroid and FWHM parameters as described in 3.3. Both polarizations are measured independently in subsequent transits at the same zenith angle. Figure 8 shows the best-fit centroid values from the drone and holography data and compares them to photographic measurements performed along the CHIME feedline to measure feed displacement along the focal line [46] Because the solar data does not include any feed-level data, it is not represented in Figure 8. Although the drone measurements are in the near-field, the centroid values and trend along the focal line from the drone agree with the two far-field measurements and follow the feed displacements. The centroid trend is known, and was previously published in [1].
Figure 8: Best-fit centroid position from Drone and Holography data sets as a function of TelX for both $X$ (left panel) and YY (right panel) polarizations. Median centroid positions for all 256 feeds (median across all 1024 frequencies) from the drone data set are shown (navy dots). Shaded navy contours indicate 68.2% and 95.4% confidence intervals. Centroid positions from holography of Virgo A are similarly indicated (yellow dots), and can be compared with similar plots in [1, 32]. These centroid measurements are consistent with photogrammetric measurements of the feed displacement presented in a doctoral thesis[46] (large purple dots). All data sets have been mean-subtracted such that absolute offsets are not included.
Figure 9: Best-fit FWHM from Drone, Holography (both for a single feed and an average of baselines $\leq 10m$ to form a data set equivalent to the solar data), and Solar data sets as a function of TelX. These curves can be directly compared because they all originate from a TelY of -0.6, where all three data sets overlap (see Figure 4). The holography and drone measurements are most comparable, because they can be separated into individual feeds. Conversely, the solar measurements are comprised of an average of all intercylinder baselines under 10m. To examine the divergence of the YY Polarization drone FWHM curve from the solar and holography datasets, we investigate transit data in the frequency band (676.2-695.7 MHz) within the purple region in Figure 10. To examine trends in FWHM as a function of both frequency and TelY, refer to Figure 6.
The best-fit FWHM parameter at TelY $\approx-0.6$ as a function of frequency is shown in Figure 9 for three data sets: the solar data (formed from combining baselines less than 10 m), the drone and holography data from a single feed, and the holography data with feeds within 10 m combined to form a data set equivalent to the solar data. The characteristic 30 MHz ripple feature is aligned for both polarizations at all frequencies, as one would expect from Figure 6. The amplitude of the ripple feature is generally comparable between all data sets for the XX polarization, though the solar FWHM is wider than the holography and drone data sets throughout the frequency band. For the YY polarization, we find that the combined-feed holography data matches its solar counterpart very well, while the ripple amplitude is slightly higher for the single-feed holography data compared to the combined-feed holography data. This suggests that feed-to-feed variations may influence the amplitude of the ripple feature. We also find the peak FWHM amplitude found in the drone data is typically significantly higher than the other data sets for YY polarization, which prompted further investigation.
Figure 10 contains drone transit measurements for frequencies within 676.2-695.7 MHz. This frequency band follows the evolution of the 30 MHz ripple feature in the FWHM(f)(f) curve from a maximum to a minimum, indicated by purple region in Figure 9. In both polarizations, we expect to find a well defined main beam at TelX=0.0 that ends at a first null around $\pm 0.04\leq\textrm{TelX}\leq\pm 0.05$). Wherever the 30 MHz ripple feature reaches a local minimum, (e.g. $f=695$ MHz plotted in black in Figure 10,) the main beam and sidelobes are discrete and the best-fit Gaussian FWHM values agree with the solar and holography datasets as expected for both the XX (top panel) and YY (bottom panel) polarizations. Conversely, wherever the 30 MHz ripple reaches a local maximum, the primary beam and first sidelobe are not separated by a null in the drone transit data. Instead, the main beam and first sidelobe are combined and the Gaussian fit finds a FWHM that is wider than expected. Examples can be found in both the XX polarization (e.g. $f=685.0$ MHz plotted in pink in the top panel of Figure 10) and YY polarization (e.g. $f=677.5$ MHz plotted in yellow in the bottom panel of Figure 10).
To demonstrate this behavior, we compare both polarizations at $f=677.5$ MHz (plotted in yellow). The XX polarization data clearly recovers the main beam and two well-defined sidelobes. The YY polarization data only recovers the second sidelobe (centered at $\textrm{TelX}\approx\pm 0.1$) while the first sidelobe (centered at $\textrm{TelX}\approx\pm 0.05$ in XX) been been incorporated into the main beam. This behavior evolves smoothly in both polarizations as we progress along the gradient of plotted frequencies. The XX polarization recovers the expected first null on either side of the main beam at the highest and lowest plotted frequencies, but intermediate frequencies show the first sidelobe joining the main beam (particularly for $\textrm{TelX}<0$). The evolution for the YY polarization is similar but not identical—as the plotted frequency decreases, the sharpness of the null diminishes, eroding the distinction between the main beam and the first sidelobe.
Figure 10: First eigenmode data from autocorrelation measurements of the drone transiting at TelY $\approx-0.6$ (feed 210) across CHIME cylinder C. The data shown are not normalized or corrected with the best-fit centroid. The frequency band shown here is reversed from the frequency axis of Figure 9. The colored curves thus evolve along the 30 MHz ripple feature in the FWHM(f)(f) curve starting from a minima at 695.7 MHz (black), where the FWHM(f)(f) values are in agreement with the solar and holography datasets, to a maxima at 676.2 MHz (yellow), where the divergence is highest. In the logarithmic scale shown in Figure 10 the amplitude appears compressed, but in linear units the amplitude does not appear compressed.
4.2 Trends with Frequency and TelY
The solar and drone data can be compared in the TelX vs. frequency and TelX vs. TelY subspaces to further investigate beam patterns. The drone data is taken from the first eigenmode (see Section 2.2.3) to increase the signal to noise. As previously described, the solar data is an average of all intercylinder baselines $\leq 10m$. All data has been corrected for the frequency-dependent centroid deviation measured in the Gaussian fits.
First, we compare the drone and solar data as a function of TelX and frequency in Figure 11. For this comparison we select a single TelY index for the solar transit data and a single CHIME feed index for the drone data (both corresponding to TelY=0.600TelY=-0.600, feed 210). The amplitude of the drone and solar data have been normalized in each frequency with the amplitude of the associated best-fit gaussian. These waterfall plots show similar frequency dependent variations in width in the main beam and sidelobe amplitudes and positions. Specifically, the width decreases with increasing frequency, and is modulated by a 30 MHz ripple aligned between the two data sets (as would be expected from Figure 9). Similarly, the XX polarization has a wider overall beam than the YY polarization, also as expected from the CHIME feed illumination pattern[47, 1]. The drone data has a more pronounced ripple in the XX polarization, whereas the ripple is attenuated in the solar data due to averaging over many baselines (supported by the lower ripple amplitude as described in Section 2.3.3).
For both polarizations, the main beam and sidelobes have similar TelY dependence in both the drone and solar datasets. A notable difference between these datasets is the high signal-to-noise and continuity present in the solar data from coadding in baseline and time. Conversely, each of the TelY values in the drone data comes from the same 3.4 min drone flight, and does not benefit from coadding or averaging. However, signal-to-noise improvements are possible with modest increases in flight time.
@@ 4.2 Trends with Frequency and TelY
Turning to the XX polarization (top panels of Figure 12), we observe the same similarities across the drone and solar data. The main beam broadens at TelY=-0.55, and shrinks to a waist at TelY=-0.52 and TelY=-0.59 in both data sets. Unlike in polarization YY, the sidelobe levels for polarization XX evolve in TelX and TelY simultaneously—shown by symmetric sloped contours that extend away from waist in the main beam (e.g. starting at TelY=-0.59, and expanding in TelX as TelY decreases to -0.62.)
@@
Figure 11: Waterfall plots of telescope data obtained during Drone measurements (Right Column, single feed [210] in either polarization) and Solar measurements (Left Column, average of intercylinder baselines $\leq 10m$ in either polarization). Similar widths and features are present in both data sets, and the beamwidths and sidelobe features are comparable. Each frequency is normalized independently to a Gaussian fit peak value to remove the uncalibrated frequency dependence of the drone transmitter.
@@
Figure 12: Waterfall plots of telescope data obtained during Solar measurements (average of intercylinder baselines $\leq 10$ for a single frequency bin (center 690MHz) in $X$ (Left Top) and YY (Left Bottom) polarizations) and Drone measurements (all feeds for a single frequency bin (center 690 MHz) in $X$ (Right Top) and YY (Right Bottom) polarizations).